Crane Company purchased equipment in January of 2008 for $390000. The equipment was being depreciated on the straight-line method over an estimated useful life of 20 years, with no salvage value. At the beginning of 2018, when the equipment had been in use for 10 years, the company paid $48000 to overhaul the equipment. As a result of this improvement, the company estimated that the useful life of the equipment would be extended an additional 5 years. What should be the depreciation expense recorded for this equipment in 2018? $16200 $13050 $9750 $19500
Answer #1

Depreciation per year = (Cost-Scrap)/Salvage value = (390000-0)/20 = 19500

The equipment has been use for 10 years, therefore total depreciation till 2018 = 19500*10 = 195000

WDV of equipment after 10 years = 395000-195000= 195000

At the beginning of 2018, revised WDV = 195000+48000 = 243000

Useful life = Balance 10 years remaining + additional 5 years = 15 years

Depreciation for 2018 = 243000/15 = $16200